OpenStudy (anonymous):

I need to know how to figure out the following: How much to invest now at 6.25% compounded daily to have $100,000 in 17 years.

OpenStudy (phi):

do you know the formula for future value?

OpenStudy (anonymous):

no

OpenStudy (btaylor):

\[A=P(1+\frac{r}{n})^{nt}\] \[100000=P(1+\frac{.0625}{365})^{365 \times 17}\] \[100000=P(1.00017123)^{6205}\] \[100000=P(2.893)\] \[P=$34,562.84\]

OpenStudy (anonymous):

thank you very much :)